OpenStudy (anonymous):

3y+x>2 and x+2y<=3 what can be said about the value of y?

OpenStudy (anonymous):

It is a value

OpenStudy (anonymous):

Just joking

OpenStudy (anonymous):

Solve for y in both equations

OpenStudy (anonymous):

3y + x > 2 3y > 2 - x y > (2 - x) / 3 x + 2y <= 3 2y <= 3 - x y <= (3 - x) / 2 (2 - x) / 3 < y <= (3 - x) / 2